OpenStudy (anonymous):

2cos theta = sec theta.. what are they asking me to find?

OpenStudy (anonymous):

\[\theta\]

OpenStudy (anonymous):

\[2\cos\theta=\sec\theta\]

OpenStudy (anonymous):

the exact values of the degree?

OpenStudy (anonymous):

i would think so

OpenStudy (anonymous):

on what interval?

OpenStudy (anonymous):

0 degree to 360 degree

OpenStudy (anonymous):

\[2\cos{\theta}=\sec{\theta} \rightarrow 2\cos{\theta}=\frac{1}{\cos{\theta}}\]then\[2\cos^2{\theta}=1\]

OpenStudy (anonymous):

\[\cos{\theta}={\pm}\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

Find all those angles theta that satisfy this.

OpenStudy (anonymous):

θ = 45, 215, 135, 225

OpenStudy (anonymous):

itsmesowat.. how'd you get the answers?

OpenStudy (anonymous):

use your unit circle or other favorite method to get special angle values for consine...

OpenStudy (anonymous):

use the calculator do [2nd] [cos] of the fraction 1/rad(2)

OpenStudy (anonymous):

Draw a right-angled triangle with sides 1, 1 and hypotenuse sqrt(2). cos(theta)=1/sqrt(2) when theta = 45 degrees. In the second quadrant, 180-45 = 135 degrees, in the third, 180+45 degrees = 225 and in the fourth, 360-45 = 315 degrees/

OpenStudy (anonymous):

gotcha.

OpenStudy (anonymous):

you got everything lokisan said? you are a genius!!~~~~

OpenStudy (anonymous):

lol tyler.. where'd you get 215 degrees?

OpenStudy (anonymous):

215 is typo, should be 225, which is 180+45, reference angle of 135

OpenStudy (anonymous):

gotcha.

OpenStudy (anonymous):

sorry i suck at typing out math... they should introduce the white board on this haha

OpenStudy (anonymous):

lmao. XD

OpenStudy (anonymous):

they have these equation thingys. it's built from mathjax

OpenStudy (anonymous):

wait. how'd you get 1/sqrt of 2?

OpenStudy (anonymous):

2cos^2(theta)=1. Divide both sides by 2, then take the sqrt of both sides.